# Dark Matter and General Relativistic Instability in Supermassive Stars

### v2

@@ Abstract
We calculate the extent to which collisionless dark matter impacts the stability of supermassive stars $(M\gtrsim 10^{4}\,M_{\odot})$. We find that, depending on the star’s mass, a dark matter content in excess of ${\sim}1\%$ by mass throughout the entire star can raise the critical central density for the onset general relativistic instability, in some cases by orders of magnitude. We consider implications of this effect for the onset of nuclear burning and significant neutrino energy losses.
@@
[…]
[lines=5]In this paper, we examine how dark matter (DM) could impact the onset of general relativistic instability in massive self-gravitating configurations, i.e., supermassive stars (SMSs), as well as their nuclear and neutrino evolution. SMSs are supported principally by radiation pressure, making them vulnerable to general relativistic instability. If these extreme objects had existed in the early universe, and then collapsed to black holes, they could provide massive seeds that facilitate early production of supermassive black holes (SMBHs).
@@
In Section II, we examine two kinematic limits of the DM based on its specific kinetic energy relative to the gravitational potential at the SMS’s surface, determine the total energy added by this DM, and derive how its presence impacts the critical central density as a function of the DM mass fraction. We then present our results for several SMS masses to examine how the effect also depends on the total mass.
In Section III, we approximate how the total fraction of DM changes with radial perturbations of the SMS, providing insight into the physical mechanism for the enhanced stability discussed in the previous section.
Finally, in Section IV, we investigate nuclear burning rates, neutrino emission rates, and their respective luminosities as a fraction of the stars’ Eddington luminosities. This allows us to ascertain how “important” nuclear burning is to an SMS’s evolution and when neutrino emission overpowers it. We also determine by how much these luminosities change with the addition of DM, and what that may mean for the downstream stability and entropy evolution.
I Stability of Supermassive Stars
@@
where $\mu=0.59$ is the mean molecular weight (amu) per particle for a primordial composition gas.
@@
where $P$ is the total pressure, $s$ is the entropy per baryon, $\rho$ is the mass density, and $\Gamma$ is the so-called polytropic exponent, with the associated quantity $n$ being the polytropic index. For a radiation dominated system, we can safely assume that the radiation field dominates the entropy per baryon $s\approx s_{\text{rad.}}=\tfrac{4}{3}aT^{3}/n_{\text{b}}$, where $T$ is the plasma temperature, $n_{\text{b}}=\rho/(\mu m_{\text{b}})$ is the baryon number density and $m_{\text{b}}\approx 931.5$ MeV is the mass of an atomic mass unit. Here, in natural units, $a=\pi^{2}/15$ is the radiation constant. Since $s$ is constant throughout the fully convective SMS, we can use Eq. (1) and $P_{\text{gas}}=n_{\text{b}}T$ and $P_{\text{rad.}}=aT^{4}/3$ to derive
$s\approx\frac{4\mu}{4.3}\left(\frac{M}{M_{\odot}}\right)^{1/2}-4,$ (3)
@@
II Collisionless Dark Matter
@@
The final term in the above equation represents the ratio of the gravitational potential of a DM particle to its specific kinetic energy. It permits two natural limits: $\tfrac{1}{2}v_{\infty}^{2}\gg GM(r)/r$ and $\tfrac{1}{2}v_{\infty}^{2}\ll GM(r)/r$ which, following Ref. [24]’s convention, we will call “hot” and “cold” DM, respectively. It should be noted that this nomenclature is not to be confused with the cosmological notion of “Hot Dark Matter” and “Cold Dark Matter,” which refers to the kinematics of DM at its decoupling epoch and how these kinematics affect early structure formation. Here, we will assume that both our “hot” and “cold” limits refer to DM particles with non-relativistic kinematics.
II.1 “Hot” Dark Matter
@@
$=\frac{8\pi^{2}}{3}G\rho_{\text{DM}}\int_{0}^{R}\rho(r)r^{2}\,\mathrm{d}r+CM.$ (17)
We will assume that $M$, $\mathrm{d}m$, and $\rho(r)$ are dominated by the baryon rest mass. Note that $\rho(r)$ very closely follows the profile of an $n=3$ polytrope. For such a polytrope in Newtonian gravitation, $M$ is independent of the central density, so we will ignore it for the purposes of extremization with respect to the central density $\rho_{\text{c}}$. The integral can be evaluated by employing the well known solutions to the Lane-Emden Equation of order 3, yielding
@@
$\tilde{E}=k_{1}\tilde{M}^{1/3-\beta/3}x^{1+\beta/2}-k_{2}\tilde{M}x-k_{4}\tilde{M}x^{2}\\ +k_{\text{HDM}}\tilde{M}^{1/3}x_{\text{DM}}^{3}x^{-2}.$ (19)
Extremization of this yields the following set of equations to be solved simultaneously after expanding in powers of the ratio of gas pressure to total pressure $\beta$:
@@
II.2 “Cold” Dark Matter
@@
where we have also made the same approximation as Butler et al. [24] that since an $n=3$ polytrope is so centrally condensed, it is reasonable to treat the SMS as a point source of mass $M$. Without this assumption, the DM density profile would feature a flattened core as opposed to the above divergent cusp at the origin. Such a flat shape would be closer to the limit described in Section II.1, so the ‘cuspy’ profile offers sufficiently disparate extreme from the “hot” DM case to allow us to explore the full range of DM effects we would expect to see.
@@ III DM Response to SMS Contraction
III.1 “Hot” Dark Matter Case
@@
III.2 “Cold” Dark Matter Case
@@
III.3 Physical Origin of Enhanced Stability
The stability of a purely standard model star (i.e., composed of baryons, electrons, photons, etc.) is determined by its pressure-averaged adiabatic index as given in Eq. (5). A heuristic physical picture for this effective “spring constant” for the star’s material is as follows: Squeeze the star radially and rapidly (so rapidly that heat does not flow and so entropy remains constant) to higher density and ask how much the pressure increases – i.e., how much does the material “push back” when quickly perturbed to higher density? Of course, when the pressure support for a self-gravitating configuration stems from particles with relativistic kinematics (e.g., photons in the SMS case), that configuration is close to instability. For example, this case would correspond to a zero total energy configuration in purely Newtonian gravitation, therefore pressure forces will grow in lock-step with the increased gravitational force from compressing the star. This means the configuration is in a meta-stable state as it costs no energy to expand or contract the star, making it ripe for instability once the non-linearity of GR comes into play.
@@ III.3 Physical Origin of Enhanced Stability
IV Nuclear Burning and Neutrino Losses
@@
It should be expected that, at least for some masses in the range ${\sim}10^{4}\text{--}10^{8}\,M_{\odot}$, nuclear (hydrogen) burning on the PPI chain is taking place at a high enough rate to significantly contribute to the total energy budget of the SMS. This rate depends on both the density and temperature throughout the star, both of which decrease as the mass increases. Specifically, at the onset of collapse where the density and temperature are highest, $\rho_{\text{c}}\sim M^{-7/2}$ and $T_{c}\sim M^{-1}$, the latter following from the definition of an $n=3$ polytrope. Since including DM can raise the critical central density (and hence the temperature) required before collapse, we can also investigate how the DM mass fraction impacts the energy released by the PPI chain throughout the star. Chen et al. [31] investigates Helium burning in SMSs. This could be relevant for SMSs light enough to have a stable Hydrogen burning phase.
@@
IV.1 PPI Chain Nuclear Burning
@@ V Conclusion
In Section I, we derived the standard result for the critical central density beyond which a hydrostatic SMS is unstable to collapse. This rested upon extremizing the total energy of the star with respect to the central density. We found that, as expected, higher mean molecular weight plasmas and higher mass SMSs are more unstable to collapse. Both of these factors serve to decrease the critical central density.
@@
In Sections II.1 and II.2, we derived analytically how the critical central density changes as a function of the DM mass fraction within the star. We considered two limiting cases of the DM kinematics, the “hot” case and the “cold” case, which compared the typical DM particle’s specific kinetic energy far away from the SMS to the gravitational potential felt at the SMS surface, and found that in both cases the critical central density for the onset of general relativistic instability increased as the DM mass fraction increases. The “cold” case shows a slightly stronger effect, and the relative increase is greater for more massive stars in both kinematic limits.
@@
In Section III we investigated a possible mechanism for the increased stability. We showed that if the star contracts, the total mass fraction of DM inside the star decreases, so the total mass enclosed by the star also decreases. This added stability is therefore completely agnostic to any particle physics specific to the DM, so long as any consideration of such physics keeps the DM non-relativistic and non-interacting with the baryonic matter. In fact, if the DM possesses some self-annihilation cross-section, like in the model discussed in Freese et al. [34], this may further bolster the stability by providing an extra source of energy and entropy.
@@
Finally, in Section IV, we considered by what degree Hydrogen burning could impact the evolution of these SMSs and how DM influences that burning. We also analyzed how neutrino emission from the high-entropy plasma would take away energy, potentially faster than what is generated from fusion, as the star approaches the critical central density from below. We found that, considering both energy generation from fusion and energy losses from neutrino emission, DM could allow heavier stars than previously predicted to stably burn on the main sequence at some point in their lives, increasing the rough upper limit of ${\sim}30,000\,M_{\odot}$ to about $50,000\,M_{\odot}$, depending on when one considers nuclear burning to be “important” to the evolution of the star.
@@
Here we have used a very simplistic model, a non-rotating and isolated SMS, to explore the physical effects of a large DM content. Nonetheless, our findings pose several open questions. First, with the inclusion of DM, what is the interplay of nuclear burning, hydrodynamics, energy transport, neutrino emission, etc., in the evolution of the SMS up to instability and the subsequent collapse to a black hole? Second, are there specific observational “handles” on the collapse of a SMS with or without DM? For example, is the gravitational radiation signature [12] different in these two cases, and to what extent would it depend on the amount of DM? Third, in this paper we have not considered the effects of DM on scenarios where a supermassive core undergoes rapid accretion, as in the hylotropic configurations proposed in Ref. [44]. The stabilizing effect of DM on such hylotropic configurations has been recently considered [45]. These scenarios may provide the most plausible route for the production of an SMBH from an SMS-like object.

### v3

@@ Abstract
We calculate the extent to which collisionless dark matter impacts the stability of supermassive stars $(M\gtrsim 10^{4}\,M_{\odot})$. We find that, depending on the star’s mass, a dark matter content in excess of ${\sim}1\%$ by mass throughout the entire star can raise the critical central density for the onset of general relativistic instability, in some cases by orders of magnitude. We consider implications of this effect for the onset of nuclear burning and significant neutrino energy losses.
@@
I Introduction
In this paper, we examine how dark matter (DM) could impact the onset of general relativistic instability in massive self-gravitating configurations, i.e., supermassive stars (SMSs), as well as their nuclear and neutrino evolution. SMSs are supported principally by radiation pressure, making them vulnerable to general relativistic instability. If these extreme objects had existed in the early universe, and then collapsed to black holes, they could provide massive seeds that facilitate early production of supermassive black holes (SMBHs).
@@ I Introduction
It will be shown in what follows that the amount of DM needed to substantially change where this instability occurs is rather extreme. For example, the DM density at the galactic center or in DM halos considered in standard cosmology are small compared to the those considered here. However, the early universe and some proposed DM candidates, e.g. self interacting models, may provide the necessary DM densities in the compact configurations considered below. It has been argued that the DM density can be greatly increased via adiabatic contraction, a purely gravitational process. For example, this has been proposed in works on “Dark Stars,” e.g. in Ref. [24]. We return to these issues in the conclusion of this paper.
@@
In Section III, we examine two kinematic limits of the DM based on its specific kinetic energy relative to the gravitational potential at the SMS’s surface, determine the total energy added by this DM, and derive how its presence impacts the critical central density as a function of the DM mass fraction. We then present our results for several SMS masses to examine how the effect also depends on the total mass.
In Section IV, we approximate how the total fraction of DM changes with radial perturbations of the SMS, providing insight into the physical mechanism for the enhanced stability discussed in the previous section.
In Section V, we determine the extent to which the equilibrium SMS density distribution is impacted by explicitly considering the gravitational influence of the DM. This analysis is performed for several SMS masses and DM mass fractions.
Finally, in Section VI, we investigate nuclear burning rates, neutrino emission rates, and their respective luminosities as a fraction of the stars’ Eddington luminosities. This allows us to ascertain how “important” nuclear burning is to an SMS’s evolution and when neutrino emission overpowers it. We also determine by how much these luminosities change with the addition of DM, and what that may mean for the downstream stability and entropy evolution.
II Stability of Supermassive Stars
@@
where $\mu=0.59$ is the mean molecular weight (in atomic mass units) per particle for a primordial composition gas.
@@
where $P$ is the total pressure, $s$ is the entropy per baryon, $\rho$ is the mass density, and $\Gamma$ is the so-called polytropic exponent, with the associated quantity $n$ being the polytropic index. For a radiation dominated system, we can safely assume that the radiation field dominates the entropy per baryon $s\approx s_{\text{rad.}}=\tfrac{4}{3}m_{\mathrm{b}}aT^{3}/\rho$, where $T$ is the plasma temperature, and $m_{\text{b}}\approx 931.5$ MeV is the mass of an atomic mass unit. Here, in natural units, $a=\pi^{2}/15$ is the radiation constant. Since $s$ is constant throughout the fully convective SMS, we can use Eq. (1) and $P_{\text{gas}}=\rho T/(\mu m_{\mathrm{b}})$ and $P_{\text{rad.}}=aT^{4}/3$ to derive
$s\approx\frac{4}{4.3}\left(\frac{M}{M_{\odot}}\right)^{1/2}-\frac{4}{\mu},$ (3)
@@
III Collisionless Dark Matter
@@
The final term in the above equation represents the ratio of the gravitational potential of a DM particle to its specific kinetic energy. It permits two natural limits: $\tfrac{1}{2}v_{\infty}^{2}\gg GM(r)/r$ and $\tfrac{1}{2}v_{\infty}^{2}\ll GM(r)/r$ which, following Ref. 25’s convention, we will call “hot” and “cold” DM, respectively. It should be noted that this nomenclature is not to be confused with the cosmological notion of “Hot Dark Matter” and “Cold Dark Matter,” which refers to the kinematics of DM at its decoupling epoch and how these kinematics affect early structure formation. Here, we will assume that both our “hot” and “cold” limits refer to DM particles with non-relativistic kinematics.
III.1 “Hot” Dark Matter
@@
$=\frac{8\pi^{2}}{3}G\rho_{\text{DM}}\int_{0}^{R}\rho(r)r^{4}\,\mathrm{d}r+CM.$ (17)
We will assume that $M$, $\mathrm{d}m$, and $\rho(r)$ are dominated by the baryon rest mass. Note that $\rho(r)$ very closely follows the profile of an $n=3$ polytrope. We will extremize the total energy while keeping the mass $M$ fixed, so the final term in Eq. (17) will be ignored. The first term in Eq. (17) can be evaluated by employing the well known solutions to the Lane-Emden Equation of order 3, yielding
@@
$\tilde{E}=k_{1}\tilde{M}^{1/3-\beta/3}x^{1+\beta/2}-k_{2}\tilde{M}x-k_{4}\tilde{M}x^{2}\\ +k_{\text{HDM}}\tilde{M}^{1/3}x_{\text{DM}}^{3}x^{-2},$ (19)
where $x_{\mathrm{DM}}^{3}=G^{3}M^{2}\rho_{\mathrm{DM}}$. Extremization of this yields the following set of equations to be solved simultaneously after expanding in powers of the ratio of gas pressure to total pressure $\beta$:
@@
III.2 “Cold” Dark Matter
@@
where we have also made the same approximation as Butler et al. [25] that since an $n=3$ polytrope is so centrally condensed, it is reasonable to treat the SMS as a point source of mass $M$. Without this assumption, the DM density profile would feature a flattened core as opposed to the above divergent cusp at the origin. Such a flat shape would be closer to the limit described in Section III.1, so the ‘cuspy’ profile offers sufficiently disparate extreme from the “hot” DM case to allow us to explore the full range of DM effects we would expect to see.
@@ IV DM Response to SMS Contraction
IV.1 “Hot” Dark Matter Case
@@
IV.2 “Cold” Dark Matter Case
@@
IV.3 Physical Origin of Enhanced Stability
The stability of a purely standard model star (i.e., composed of baryons, electrons, photons, etc.) is determined by its pressure-averaged adiabatic index as given in Eq. (5). A heuristic physical picture for this effective “spring constant” for the star’s material is as follows: Squeeze the star radially and rapidly (so rapidly that heat does not flow and so entropy remains constant) to higher density and ask how much the pressure increases – i.e., how much does the material “push back” when quickly perturbed to higher density?
Of course, when the pressure support for a self-gravitating configuration stems from particles with relativistic kinematics (e.g., photons in the SMS case), that configuration is close to instability. For example, this case would correspond to a zero total energy configuration in purely Newtonian gravitation, therefore pressure forces will grow in lock-step with the increased gravitational force from compressing the star. This means the configuration is in a meta-stable state as it costs no energy to expand or contract the star, making it ripe for instability once the non-linearity of GR comes into play.
@@ IV.3 Physical Origin of Enhanced Stability
We note that the increased stability stemming from a fixed DM distribution inside our SMS may have similarity with the enhanced stability found in giant-planet envelope models with a fixed, incompressible core [32]. In both the giant-planet envelope and our SMS cases there is little or no change in the background gravitational potential of the core in the former case, or the DM envelope in the latter.
V Equilibrium Structure of Star with Dark Matter
Let us consider the “hot” dark matter case where the density profile follows a flat distribution. Considering that SMSs are well approximated by the structure of an $n=3$ polytrope, they are highly centrally condensed relative to their average density. For any non-zero DM mass fraction $f$, there should be a radius $r_{\mathrm{eq.}}<R$ beyond which the DM density is greater than the standard model matter density. Since the pressure gradient of the SMS depends on the local gravitational field strength, we should expect that around and beyond this radius the SMS structure may be altered by this extra gravitation source.
Specifically, if we assume the DM provides no direct pressure support but does interact with the SMS gravitationally, the pressure gradient at a given radius $r$ is
$\frac{\mathrm{d}P_{0}}{\mathrm{d}r}=-\frac{G(m_{0}+m_{\mathrm{DM}})(\rho_{0}+\rho_{\mathrm{DM}})}{r^{2}},$ (41)
where $P_{0}$, $m_{0}$, and $\rho_{0}$ are the standard model pressure, mass enclosed, and density, respectively, and $m_{\mathrm{DM}}$ and $\rho_{\mathrm{DM}}$ are the DM mass enclosed and density, respectively. Both $m_{0}$ and $m_{\mathrm{DM}}$ have the following continuity conditions:
$\frac{\mathrm{d}m_{0}}{\mathrm{d}r}$ $=4\pi r^{2}\rho_{0},$ (42)
$\frac{\mathrm{d}m_{\mathrm{DM}}}{\mathrm{d}r}$ $=4\pi r^{2}\rho_{\mathrm{DM}}.$ (43)
With these three equations, Eq. (41) can be manipulated to derive
$\frac{1}{r^{2}}\frac{\mathrm{d}}{\mathrm{d}r}\left(\frac{r^{2}}{\rho_{0}+\rho_{\mathrm{DM}}}\frac{\mathrm{d}P_{0}}{\mathrm{d}r}\right)=-4\pi G(\rho_{0}+\rho_{\mathrm{DM}}).$ (44)
Let us now assume that the standard model density is parameterized by its central density $\rho_{\rm 0c}$ and a function of radius $\theta$ via $\rho_{0}=\rho_{\rm 0c}\theta^{n}$, where $n$ is the polytropic index. Plugging this into Eq. (2) yields $P_{0}=K\rho_{\rm 0c}^{1+1/n}\theta^{n+1}$. Collecting these relations together in Eq. (44) gives
$\frac{1}{r^{2}}\frac{\mathrm{d}}{\mathrm{d}r}\left(\frac{r^{2}}{\theta^{n}+\frac{f}{C_{3}}}\alpha_{\mathrm{LE}}^{2}\theta^{n}\frac{\mathrm{d}\theta}{\mathrm{d}r}\right)=-\left(\theta^{n}+\frac{f}{C_{3}}\right),$ (45)
where $\alpha_{\mathrm{LE}}^{2}=(n+1)K\rho_{\rm 0c}^{1/n-1}/(4\pi G)$ is the square of the standard Lane-Emden length scale and $f/C_{3}$ is the DM mass fraction scaled by the central condensation as per Eq. (21), specifically for a flat DM distribution.
Scaling the radius $r$ via $r=\alpha_{\mathrm{LE}}\xi$, we arrive at a modified Lane-Emden equation which reads
$\frac{1}{\xi^{2}}\frac{\mathrm{d}}{\mathrm{d}\xi}\left(\frac{\xi^{2}}{\theta^{n}+\frac{f}{C_{3}}}\theta^{n}\frac{\mathrm{d}\theta}{\mathrm{d}\xi}\right)=-\left(\theta^{n}+\frac{f}{C_{3}}\right),$ (46)
which reduces to the standard Lane-Emden equation for $f\rightarrow 0$. We can solve this equation numerically for different values of $f$, as presented in Figure 2 for $n=3$. The solutions agree with the standard (no DM) case for most of the x-axis range of the figure, but begin to diverge radically at roughly where the standard model and DM densities are comparable.
Figure 2: Solutions to Eq. (46) of the function $\theta$ for different DM mass fractions $f$. The actual physical density is $\rho_{0}(r)=\rho_{\rm c0}\theta^{3}(\xi)$, where $\xi=r/\alpha_{\mathrm{LE}}$.
Using these functions, the altered values of the structure coefficients $k_{1}$, $k_{2}$, $k_{4}$, and $k_{\mathrm{HDM}}$ are tabulated as a function of DM mass fraction $f$ in Table 1. These were calculated following the prescriptions in Chapter 6 of Shapiro and Teukolsky [31]. In Figure 3, the relative change in the critical central density for several SMS masses and DM mass fractions is visualized. For all SMS masses and DM mass fractions, including the effect of the altered equilibrium distribution of the standard model matter lowers the critical central density slightly from Eq. (23).
$n=3$ $f=10^{-4}$ $f=10^{-3}$ $f=10^{-2}$ $f=10^{-1}$
$k_{1}$ 1.75579 1.75595 1.75714 1.76558 1.80913
$k_{2}$ 0.63900 0.63911 0.63999 0.64650 0.68533
$k_{4}$ 0.91829 0.91853 0.92044 0.93485 1.02386
$k_{\mathrm{HDM}}$ 3.58450 3.58409 3.58044 3.55002 3.37575
Table 1: Tabulated structure coefficients determined upon integrating the solutions to Eq. (46). The “$n=3$” column corresponds to the standard Lane-Emden solution (i.e. $f=0$), and the following columns are the values calculated for several chosen and increasing DM mass fractions $f$.
Figure 3: As a function of the DM mass fraction $f$, the relative change in the critical central density is plotted. Here, the quantity $\rho_{\mathrm{c,\,crit.NEW}}/\rho_{\mathrm{c,\,crit.OLD}}$ is the ratio of the critical central densities considering the change in the equilibrium density profile of the SMS vs. not considering such a change.
Clearly, considering the change in the SMS equilibrium density profile only confers significant changes in the critical central density for DM mass fractions $f$ in the larger end of the range we are considering (${\gtrsim}1\%$).
VI Nuclear Burning and Neutrino Losses
@@
It should be expected that, at least for some masses in the range ${\sim}10^{4}\text{--}10^{8}\,M_{\odot}$, nuclear (hydrogen) burning on the PPI chain is taking place at a high enough rate to significantly contribute to the total energy budget of the SMS. This rate depends on both the density and temperature throughout the star, both of which decrease as the mass increases. Specifically, at the onset of collapse where the density and temperature are highest, $\rho_{\text{c}}\sim M^{-7/2}$ and $T_{c}\sim M^{-1}$, the latter following from the definition of an $n=3$ polytrope. Since including DM can raise the critical central density (and hence the temperature) required before collapse, we can also investigate how the DM mass fraction impacts the energy released by the PPI chain throughout the star. Chen et al. [33] investigates Helium burning in SMSs, which could be relevant for SMSs light enough to have a stable Hydrogen burning phase. The following calculations in this section are determined for the non-altered SMS density distribution, contra the previous section’s treatment, for simplicity.
@@
VI.1 PPI Chain Nuclear Burning
@@ VII Conclusion
In Section II, we derived the standard result for the critical central density beyond which a hydrostatic SMS is unstable to collapse. This rested upon extremizing the total energy of the star with respect to the central density. We found that, as expected, higher mean molecular weight plasmas and higher mass SMSs are more unstable to collapse. Both of these factors serve to decrease the critical central density.
@@
In Sections III.1 and III.2, we derived analytically how the critical central density changes as a function of the DM mass fraction within the star. We considered two limiting cases of the DM kinematics, the “hot” case and the “cold” case, which compared the typical DM particle’s specific kinetic energy far away from the SMS to the gravitational potential felt at the SMS surface, and found that in both cases the critical central density for the onset of general relativistic instability increased as the DM mass fraction increases. The “cold” case shows a slightly stronger effect, and the relative increase is greater for more massive stars in both kinematic limits.
@@
In Section IV we investigated a possible mechanism for the increased stability. We showed that if the star contracts, the total mass fraction of DM inside the star decreases, so the total mass enclosed by the star also decreases. This added stability is therefore completely agnostic to any particle physics specific to the DM, so long as any consideration of such physics keeps the DM non-relativistic and non-interacting with the baryonic matter. In fact, if the DM possesses some self-annihilation cross-section, like in the model discussed in Freese et al. [36], this may further bolster the stability by providing an extra source of energy and entropy.
This brings up a potential limitation of our stability analysis. Our procedure was preformed under the assumption that the mass, total entropy, and angular momentum were fixed. If we include additional physics, such as rotation, or DM self interactions that inject entropy into the star, our analysis may fall short. Specifically, in Lai et al. [37], it is shown that a sufficient but not necessary condition for the onset of instability is that both the first and second derivatives of the total energy vanish for the case that total energies parameterized by multiple quantities. However, our analysis involves a sequence of configurations with only one parameter, the central density $\rho_{c}$, so this condition is sufficient and necessary.
In Section V, we have determined the equilibrium stellar density profile for an SMS embedded in a field of DM at a constant density. We note that the profile closely follows that of a standard $n=3$ polytrope until the two densities become comparable, whereupon they radically diverge to form a slightly smaller total configuration. The standard and new profiles are quite similar for low DM mass fractions, as expected. Using these new profiles, the change in critical central density is determined quantitatively by integrating them to calculate the structure coefficients that directly determine the critical central density. We find that for all SMS masses and DM mass fractions considered, considering the new equilibrium density profile would serve to slightly lower the increased critical central density due to the DM presence alone.
@@
Finally, in Section VI, we considered by what degree Hydrogen burning could impact the evolution of these SMSs and how DM influences that burning. We also analyzed how neutrino emission from the high-entropy plasma would take away energy, potentially faster than what is generated from fusion, as the star approaches the critical central density from below. We found that, considering both energy generation from fusion and energy losses from neutrino emission, DM could allow heavier stars than previously predicted to stably burn on the main sequence at some point in their lives, increasing the rough upper limit of ${\sim}30,000\,M_{\odot}$ to about $50,000\,M_{\odot}$, depending on when one considers nuclear burning to be “important” to the evolution of the star.
@@
Here we have used a very simplistic model, a non-rotating and isolated SMS, to explore the physical effects of a large DM content. Nonetheless, our findings pose several open questions. First, with the inclusion of DM, what is the interplay of nuclear burning, hydrodynamics, energy transport, neutrino emission, etc., in the evolution of the SMS up to instability and the subsequent collapse to a black hole? Second, are there specific observational “handles” on the collapse of a SMS with or without DM? For example, is the gravitational radiation signature [12] different in these two cases, and to what extent would it depend on the amount of DM? Third, in this paper we have not considered the effects of DM on scenarios where a supermassive core undergoes rapid accretion, as in the hylotropic configurations proposed in Ref. 46. The stabilizing effect of DM on such hylotropic configurations has been recently considered [47]. These scenarios may provide the most plausible route for the production of an SMBH from an SMS-like object.